OpenStudy (anonymous):

the right circular cylinder has radius r and height h. not including its bases, the surface area of the cylinder is 2(pi)rh. if this surface area not including the bases is 70(pi), what is the volume of the cylinder in terms of r?

OpenStudy (anonymous):

Do you know the equation for the volume of a cylinder?

OpenStudy (anonymous):

\[V = \pi r^2 h\] \[A_s = 2 \pi r h\] \[r = \frac {A_s}{2 \pi h}\] plug n chug

OpenStudy (anonymous):

errr, I guess it wants you to sub in h, not r. My bad. \[h = \frac{A_s}{2 \pi r}\]

OpenStudy (anonymous):

yeah it didn't really give me anything just that it a PSAT question

OpenStudy (anonymous):

Can you solve it from here?

OpenStudy (anonymous):

yeah I think so

OpenStudy (anonymous):

^^ Happy PSAT studying

OpenStudy (anonymous):

If you want to remember the volume, an easy way to see that it's just a bunch of circular areas stacked on top of each other. \[A_{circle} = \pi r^2\] Volume of right cylinder height h \[V = h (\pi r^2)\]